### v3

# On the Structure of Information

Abstract
We characterize information as risk reduction between knowledge states represented by partitions of the underlying probability space. Entropy corresponds to risk reduction from no (or partial) knowledge to full knowledge about a random variable, while information corresponds to risk reduction from no (or partial) knowledge to partial knowledge. This applies to any information measure that is based on expected loss minimization, such as Bregman information, with Shannon information and variance as prominent examples. In each case, fundamental properties like the chain rule, non-negativity, and the relationship between information and divergence are preserved. Because partitions form a lattice under refinement, our general treatment reveals how information can be decomposed into redundant, unique, and synergistic contributions, a question important in applications from neuroscience to machine learning, yet one for which existing formulations lack consensus regarding foundational definitions and can violate basic properties such as the chain rule or non-negativity. Redundancy corresponds to Aumann’s common knowledge, synergy to the gap between separately and jointly observed sources, and unique information is necessarily path-dependent, taking different values depending on what is already known. The resulting partial information decomposition is grounded directly in probability theory, avoids treating scalar information quantities as primitive compositional objects, yields non-negative terms by construction, and offers a more fine-grained credit assignment.
I Introduction
Observations matter because they change what we know. Their significance rarely lies in the raw values recorded but in what those values imply about other variables of interest, such as unobserved states of the world, future outcomes, or quantities we aim to predict. For example, the temperature shown by a living-room thermometer is not meaningful in itself, but through what it reveals about comfort, heating decisions, or related physical variables such as humidity or pressure. The purpose of an information measure is to assign a numerical value to such changes in knowledge, capturing how an observation improves our ability to learn, predict, and decide.
This viewpoint aligns with the classical foundations of information theory introduced by Shannon [72], in which information is defined as the minimal expected length of a bit string required to encode a given message. More precisely, the total information, or uncertainty, of a discrete signal $X$ can be quantified by $H_{S}(X)=\mathbb{E}[-\log\mathbb{P}_{X}(X)]$, known as the Shannon entropy of $X$. Shannon [73] and others [27] showed that his entropy plays a central role in coding theorems, thereby advancing Shannon’s original goal of characterizing the fundamental limits of communication. Specifically, $H_{S}(X)$ determines the minimum achievable average codeword length for encoding $X$ using binary strings [41]. Mutual information, $I_{S}(X;Y)=H_{S}(X)-H_{S}(X|Y)$, also admits interpretations in communication complexity, for instance as the minimum message length required to simulate conditional distributions $\mathbb{P}_{X|Y=y}$ given $\mathbb{P}_{X}$ [38].
Although Shannon originally conceived information theory in the context of channel coding and communication, its impact has extended to many other disciplines, including statistical mechanics [44], complexity theory [47], economics [75], decision theory [77], and machine learning [51]. Over the years, numerous alternative formulations of uncertainty and information measures have been developed for situations in which the length of binary strings transmitted over a channel is not the most appropriate quantifier. Examples include generalized entropies, such as Rényi entropy [62], Tsallis entropy [78], and Csiszár’s f-divergences [19], as well as structurally different approaches such as majorization, which formalizes uncertainty in terms of a non-total preorder on probability distributions [53, 1], with applications in economics [79], physics [68], quantum information [59], and decision theory [32].
On the other hand, in classical probability theory, it is common to interpret the partition $\pi(X)=\{X^{-1}(x)\}_{x}$ of the underlying sample space $\Omega$, or alternatively the associated $\sigma$-algebra $\sigma(X)$, as the information carried by a random variable $X$ [2, 28, 9]. Partitions represent the knowledge gained through measurements, say $X=x$, by specifying what is learned about the actual outcome $\omega\in\Omega$ from the value $x$, namely, that $\omega$ belongs to the block $X^{-1}(x)$ corresponding to $x$. The partition $\pi(X)$ thus determines the granularity at which subsets of $\Omega$ can be distinguished by observing $X$, i.e., observations with finer partitions are more informative than those with coarser partitions. Since not all partitions of $\Omega$ can be compared with respect to coarseness, this notion of information is only a partial order. An actual information measure, however, requires an additional objective with respect to which arbitrary partitions can be evaluated, effectively turning the partial order into a total preorder.
In this work, we focus on one family of such measures that unifies the coding-theoretic and the partition-based notions of information. Specifically, partitions are evaluated by the extent to which they enable optimization of a given objective function, when decisions are restricted to the coarse-graining level induced by the partition. We identify information in the general case as the reduction of risk when passing from one partition to another, with entropy appearing as the special case corresponding to risk reduction from the trivial partition (no knowledge) to the partition generated by the variable of interest (full knowledge). This construction subsumes a variety of known uncertainty and information measures, where Shannon entropy and mutual information, in particular, are recovered when the objective function corresponds to minimal expected codeword length. Nevertheless, many characteristic properties usually associated with Shannon’s measures, such as the chain rule and connections to divergence measures, are retained in the general case. Furthermore, the resulting linear structure reveals how total information decomposes into additive subterms, each representing risk reduction between successive levels of partial knowledge. This perspective links classical concepts from probability theory, such as Aumann’s notion of common knowledge [2], to partial information decomposition (PID) approaches [83, 33, 7], thereby grounding PID within the established framework of probability theory.
II From knowledge to information
II-A The knowledge lattice of partitions
In probability theory [45], observations, or measurements, are modeled as random variables, or signals, $X:\Omega\to\mathcal{X}$. That is, they are modeled as functions from the underlying sample space $\Omega$ to some well-behaved space $\mathcal{X}$ such as the real numbers (see Section II-B for a precise definition). The idea is that the set $\Omega$ is rich enough to jointly model the uncertainty of all measurements we are interested in. Once $\omega\in\Omega$ is sampled, the measurement $X(\omega)$ is uniquely specified.
In all but the simplest cases, the reverse is not true: observing a particular measurement $X=x$ does not generally allow us to determine a unique elementary event $\omega\in\Omega$. Instead, it only narrows down the possible $\omega$ that are compatible with this observation. Specifically, when observing $X=x$, we know $\omega$ is somewhere in the preimage
$\pi_{x}\coloneqq X^{-1}(\{x\})=\{\omega\in\Omega\mid X(\omega)=x\},$
which is also simply denoted by $\{X\,{=}\,x\}$. Since $X$ is a (deterministic) function on $\Omega$, the collection $\pi(X)\coloneqq\{\pi_{x}\}_{x}$ of non-empty preimages is a partition of $\Omega$, i.e., $\Omega=\bigcup_{x}\pi_{x}$ and $\pi_{x}\cap\pi_{x^{\prime}}=\emptyset$ if $x\neq x^{\prime}$.
Partitions encode a notion of knowledge: coarse partitions correspond to less informative measurements, finer partitions to more informative measurements. Formally, this idea is captured by the partial order (a reflexive, transitive, and antisymmetric relation) of partitions, given by $\pi\preceq\pi^{\prime}$ if and only if $\pi^{\prime}$ is finer than $\pi$, that is, if any element of $\pi$ can be written as a union of elements of $\pi^{\prime}$. Note that, in the set theory literature, this relation is sometimes defined in reverse order. We adopt the usual order-theoretic language, e.g., $\pi$ and $\pi^{\prime}$ are called comparable if $\pi\preceq\pi^{\prime}$ or $\pi^{\prime}\preceq\pi$. Even if two partitions $\pi$ and $\pi^{\prime}$ are not comparable, we can form their least upper bound $\pi\vee\pi^{\prime}$ (their join), which is their coarsest common refinement, while their greatest lower bound $\pi\wedge\pi^{\prime}$ (their meet) is their finest common coarsening.
Given a sample space $\Omega$ and a measurement with partition $\pi$ of $\Omega$, we call a set $B\subseteq\Omega$ knowable by observing $\pi$ (or simply $\pi$-knowable) if it has the property that each block of $\pi$ is either contained in $B$ or in $B^{c}$. In the measure theory literature, such sets are also known as $\pi$-sets [12], events with respect to $\pi$, or informed sets [40]. Intuitively, knowable sets $B$ do not overlap the blocks of $\pi$, they are aligned with the partition. Sets $B\subseteq\Omega$ that are knowable by each of $\pi$ and $\pi^{\prime}$ are called sets of common knowledge [2]. We call a set of common knowledge minimal if it contains no proper nonempty subset that is also a set of common knowledge.
Figure 1: Examples of partitions $\pi$ and $\pi^{\prime}$ of $\Omega=[0,1]$, including their coarsest common refinement (join $\pi\vee\pi^{\prime}$) and finest common coarsening (meet $\pi\wedge\pi^{\prime}$). The blocks of $\pi\vee\pi^{\prime}$ are knowable from the combination of $\pi$ and $\pi^{\prime}$, while the blocks of $\pi\wedge\pi^{\prime}$ are sets of common knowledge, because these sets are knowable by both $\pi$ and $\pi^{\prime}$. In the terminology of Aumann [2], a decision-maker who performs measurement $\pi$ knows the block of $\pi\wedge\pi^{\prime}$ that contains $\omega$ and also knows that a decision-maker with measurement $\pi^{\prime}$ knows that $\omega$ is in that block. They also know that the other decision-maker knows that they know, and so on.
This terminology allows for a concise description of the partial order $\preceq$ and its meet and join, which we summarize in the following definition.
Definition 1 (Knowledge lattice I).
We call a partition $\pi$ of a probability space $\Omega$ a measurement or knowledge state. The set $\Pi_{\Omega}$ of all partitions $\pi$ of $\Omega$ is partially ordered. We have:
  1. $(i)$
    $\pi\preceq\pi^{\prime}$ if and only if all observations $A\in\pi$ are knowable by observing $\pi^{\prime}$.
  2. $(ii)$
    The coarsest common refinement $\pi\vee\pi^{\prime}$ contains all non-empty intersections of elements from $\pi$ and $\pi^{\prime}$. This partition corresponds to the obtainable knowledge when measurements of both $\pi$ and $\pi^{\prime}$ are observed.
  3. $(iii)$
    The finest common coarsening $\pi\wedge\pi^{\prime}$ is the partition formed by all minimal sets of common knowledge. It corresponds to the shared knowledge obtainable by either measurement, $\pi$ or $\pi^{\prime}$.
We refer to the tuple $(\Pi_{\Omega},\preceq)$ as the knowledge lattice of partitions (of $\Omega$).
Figure 2: A small section of the knowledge lattice of a 66-element sample space $\Omega=\{1,\dots,6\}$. Partitions connected through a path are comparable with respect to $\preceq$, with upper ones being finer than lower ones. Note that, $\pi$ and $\pi^{\prime}$ are not comparable with respect to $\preceq$, while both are bound by their meet and join. $\pi$ could for example be the partition corresponding to the random variable $X$ given by $X(\omega)=\lfloor(\omega-1)/2\rfloor$ (integer division by 2 of $\omega-1$).
Partitions enable a comparison of signals based on their informational content: each signal induces a partition of the state space and finer partitions represent more informative signals. This representation has a long tradition in probability [10], decision theory [2], and economics [58]. However, continuous signals $X$ induce partitions with uncountably many blocks $\pi_{x}=\{X^{-1}(x)\}$, which are typically measure-zero sets under the underlying probability measure. To describe conditioning on such events, modern probability theory is built on $\sigma$-algebras instead, which provide the minimal structure needed to describe events and their probabilities in general spaces. To connect the intuitive picture offered by partitions to the framework of probability theory, it is therefore necessary to pass from partitions to suitable $\sigma$-algebras.
II-B The knowledge lattice of $\sigma$-algebras
A $\sigma$-algebra on $\Omega$ is a collection of subsets $A\subseteq\Omega$ that contains $\Omega$ and is closed under complements and countable unions. $\sigma$-algebras form the basic building blocks of measure theory: a set $\Omega$ together with a $\sigma$-algebra $\Sigma$ is known as a measurable space $(\Omega,\Sigma)$, elements of $\sigma$-algebras are called measurable sets, and measures $\mu$ are special mappings defined on $\sigma$-algebras that are additive under disjoint unions. Functions between measurable spaces are called $\mathcal{F}$-measurable if the pre-images of measurable sets lie in the $\sigma$-algebra $\mathcal{F}$. A measure $\mathbb{P}$ on $(\Omega,\Sigma)$ is a probability measure if $\mathbb{P}(\Omega)=1$, in which case $(\Omega,\Sigma,\mathbb{P})$ (or just $\Omega$) is called a probability space and the elements $B\in\Sigma$ are called events. A random variable (or random element) $X:(\Omega,\Sigma)\to(\mathcal{X},\mathcal{B})$ is any $\Sigma$-measurable mapping from a probability space $\Omega$ to some other measurable space $(\mathcal{X},\mathcal{B})$.
Random variables induce a sub-$\sigma$-algebra $\sigma(X)\subseteq\Sigma$, which is the smallest $\sigma$-algebra on $\Omega$ with respect to which $X$ is measurable, i.e., $\sigma(X)$ is the collection of all pre-images $X^{-1}(B)$ of measurable sets $B\in\mathcal{B}$. Unless specified otherwise, random variables are usually considered to be mappings to some Euclidean space $\mathcal{X}$, e.g., $\mathcal{X}=\mathbb{R}$, equipped with the Borel $\sigma$-algebra $\mathcal{B}(\mathcal{X})$, which is the smallest $\sigma$-algebra containing all open sets. For example, $\mathcal{B}(\mathbb{R})$ is the smallest $\sigma$-algebra that contains all open intervals $(a,b)$. The collection of sub-$\sigma$-algebras on $\Omega$ is partially ordered through inclusion and obeys a lattice structure with meet $\mathcal{F}\wedge\mathcal{F}^{\prime}$ given by the intersection of $\mathcal{F}$ and $\mathcal{F}^{\prime}$, and join $\mathcal{F}\vee\mathcal{F}^{\prime}$ given by the smallest $\sigma$-algebra containing the union of $\mathcal{F}$ and $\mathcal{F}^{\prime}$.
An important probabilistic tool is the conditional expectation $\mathbb{E}[X|\mathcal{F}]$ of a random variable $X$ with respect to the $\sigma$-algebra $\mathcal{F}$. In particular, it enables conditioning on other (even continuous) random variables $Y$ by choosing $\mathcal{F}=\sigma(Y)$. Intuitively, $\mathbb{E}[X|\mathcal{F}]$ corresponds to a coarse-graining of $X$ compatible with the knowledge encoded by $\mathcal{F}$. In fact, one of the many equivalent ways to define conditional expectation [63] is as the best such approximation in terms of mean squared error,
$\mathbb{E}[X|\mathcal{F}]=\operatorname*{\arg\!\min}_{Z\ \mathcal{F}\text{-measurable}}\mathbb{E}[(X-Z)^{2}].$ (1)
That is, $\mathbb{E}[X|\mathcal{F}]$ is the $L^{2}$-projection of $X$ onto the subspace of $\mathcal{F}$-measurable functions. It can even be shown that it is the unique $\mathcal{F}$-measurable predictor that minimizes the discrepancy to $X$ with respect to any Bregman loss, and that this property characterizes Bregman divergences [5].
Hence, the conditional expectation $\mathbb{E}[X|\mathcal{F}]$ interpolates between the single number $\mathbb{E}[X]=\mathbb{E}[X|\{\emptyset,\Omega\}]$ representing $X$ under no knowledge, and the full random variable $X=\mathbb{E}[X|\sigma(X)]$, suggesting to model knowledge through $\sigma$-algebras. The tower property further supports this interpretation, since $\mathbb{E}[\mathbb{E}[X|\mathcal{F}^{\prime}]|\mathcal{F}]=\mathbb{E}[X|\mathcal{F}]$ whenever $\mathcal{F}\subseteq\mathcal{F}^{\prime}$, i.e., under the smaller $\sigma$-algebra $\mathcal{F}$, $X$ cannot be distinguished from its coarse-grained version $\mathbb{E}[X|\mathcal{F}^{\prime}]$.
Since knowledge states and measurements were represented in the previous section as partitions $\pi$ of the sample space, these properties motivate passing to the associated $\sigma$-algebras $\sigma(\pi)$, which capture precisely the events that are distinguishable under a given state of knowledge. In particular, there are further properties of $\sigma$-algebras that align well with the intuition of informational content. For example, if $Z$ is a $\sigma(Y)$-measurable random variable, that is, if $\sigma(Z)\subseteq\sigma(Y)$, then by the Doob-Dynkin Lemma [23], $Z=f(Y)$ for some measurable $f$, i.e., if $Y$ is known then $Z$ is known.
In fact, in the discrete case, partitions $\pi$ and $\sigma$-algebras have a very simple relationship: if $\pi=\{\pi_{i}\}_{i}$ is a countable partition, for example generated by a discrete random variable $X$, where $\pi_{i}=X^{-1}(x_{i})$, then the smallest $\sigma$-algebra that contains $\pi$ is given by the collection of all $\pi$-knowable sets, i.e., all unions of blocks from $\pi$, $\sigma(\pi)=\{\bigcup_{i\in I}\pi_{i}\,|\,I\subseteq\mathbb{N}\}$. It then follows immediately that
$\pi\preceq\pi^{\prime}\quad\Leftrightarrow\quad\sigma(\pi)\subseteq\sigma(\pi^{\prime}),$ (2)
because the property that the blocks of $\pi$ are countable unions of blocks of $\pi^{\prime}$ translates directly to the corresponding $\sigma$-algebras.
However, this does not directly transfer to the continuous case, because $\sigma$-algebras are only closed with respect to countable unions, and so the smallest $\sigma$-algebra that contains an uncountable partition $\pi=\{\pi_{x}\}_{x}$ is the collection of all subsets $B\subseteq\Omega$ such that $B$ or its complement $B^{c}$ is the union of a countable number of blocks of $\pi$. This generally results in $\sigma$-algebras that are too small to satisfy (2), as can be seen from the following example from [24]: Consider the Borel space $([0,1],\mathcal{B})$ and the two partitions of $[0,1]$ given by
$\pi=\{[0,1/2),[1/2,1]\},\quad\pi^{\prime}=\{\{x\}|x\in[0,1]\}.$
Then $\pi\preceq\pi^{\prime}$, and the corresponding smallest $\sigma$-algebra that contains $\pi$ is $\mathcal{F}=\{\emptyset,[0,1/2),[1/2,1],[0,1]\}$, whereas the smallest $\sigma$-algebra that contains $\pi^{\prime}$ is the collection of subsets of $[0,1]$ that are either countable or have countable complement. In particular, this $\sigma$-algebra does not contain $\mathcal{F}$ (they are not comparable with respect to inclusion).
Rather than defining $\sigma(\pi)$ as the smallest $\sigma$-algebra containing the blocks of $\pi$, one chooses $\sigma(\pi)$ to be the $\sigma$-algebra induced by the corresponding random variable, which exists whenever $\pi$ is a measurable partition [64, 12]. Such partitions have a basis $\{B_{i}\}_{i\in I}$, that is, a countable collection of $\pi$-knowable sets that separates the blocks of $\pi$ in the sense that for any two blocks $C$ and $D$ of $\pi$, there exists $i\in I$ such that either $C\subseteq B_{i}$ and $D\not\subseteq B_{i}$, or $C\not\subseteq B_{i}$ and $D\subseteq B_{i}$. In our example above, a basis for $\pi^{\prime}$ would be the countable collection of all open sub-intervals of $[0,1]$ with rational endpoints. Another standard assumption is that the relations $\pi\preceq\pi^{\prime}$ and $\pi=\pi^{\prime}$ are considered only up to null sets, i.e., sets of measure 0 with respect to the underlying measure $\mathbb{P}$ [12]. It then follows that the meet $\pi\wedge\pi^{\prime}$ and join $\pi\vee\pi^{\prime}$ of two measurable partitions $\pi$ and $\pi^{\prime}$ are again measurable [65]. The $\sigma$-algebra $\sigma(\pi)$ of a measurable partition $\pi$ is the collection of all measurable $\pi$-sets. For our example, $\sigma(\pi^{\prime})$ is therefore the full Borel algebra $\mathcal{B}([0,1])$, and so $\sigma(\pi)\subset\sigma(\pi^{\prime})$.
The $\sigma$-algebras induced by measurable partitions are in one-to-one correspondence with the underlying partitions, and, importantly, (2) is satisfied up to null sets [65]. Moreover, the knowledge lattice of measurable partitions directly translates to their $\sigma$-algebras [65]: the common knowledge partition $\pi\,{\wedge}\,\pi^{\prime}$ is represented by the meet of their induced $\sigma$-algebras, whereas $\pi\,{\vee}\,\pi^{\prime}$ is represented by their join. These observations motivate the following definition.
Definition 2 (Knowledge lattice II).
Given a measurable partition $\pi$, the $\sigma$-algebra $\sigma(\pi)$ induced by $\pi$ is the collection of all measurable sets knowable by observing $\pi$. Then [65]:
  1. $(i)$
    $\sigma(\pi)$ represents the knowledge in $\pi$ in the set of $\sigma$-algebras on $\Omega$, i.e.,
    $\pi\preceq\pi^{\prime}\ \Leftrightarrow\ \sigma(\pi)\subseteq\sigma(\pi^{\prime})\,.$ (3)
  2. $(ii)$
    The knowable sets with respect to the finest common coarsening of two partitions $\pi$ and $\pi^{\prime}$ are exactly those sets that are knowable by both $\pi$ and $\pi^{\prime}$, i.e., the sets that are shared by $\sigma(\pi)$ and $\sigma(\pi^{\prime})$,
    $\sigma(\pi\wedge\pi^{\prime})=\sigma(\pi)\wedge\sigma(\pi^{\prime})\,.$ (4)
    Similarly, the knowable sets with respect to the coarsest common refinement of $\pi$ and $\pi^{\prime}$ form the smallest $\sigma$-algebra containing $\sigma(\pi)\cup\sigma(\pi^{\prime})$,
    $\sigma(\pi\vee\pi^{\prime})=\sigma(\pi)\vee\sigma(\pi^{\prime}).$ (5)
We refer to the set of $\sigma$-algebras induced by measurable partitions equipped with the partial order of inclusion as the knowledge lattice of $\sigma$-algebras (of $\Omega$), where meet and join operations are given by (4) and (5), respectively. The knowledge lattice of $\Omega$ can refer to either of the equivalent knowledge lattices of partitions (Definition 1) or $\sigma$-algebras.
Note that measurable partitions and their induced $\sigma$-algebras are exactly those partitions and respective $\sigma$-algebras that are generated by random variables. In fact, a classical result states that a partition is measurable if and only if its blocks can be represented as level sets $\pi_{x}=f^{-1}(\{x\})$ of some measurable function $f:\Omega\to[0,1]$ [12]. From now on, all partitions are assumed to be measurable.
II-C Risk minimization
As we have seen in the previous sections, knowledge about $\omega\in\Omega$ obtainable through measurements, represented by the underlying partitions, or equivalently by their $\sigma$-algebras, is only partially ordered. However, by explicitly specifying how useful an observation is, we can compare any partition with respect to that objective, effectively turning the knowledge lattice into a total preorder. For example, we could use the surprise generated by a particular observation about some unknown quantity, in terms of the reduction in codeword length ($\log$ loss), or the reduction in variance of an unknown signal (square error), etc. In general, information measures can be seen as risk reduction with respect to a chosen risk function $R$ defined on measurements $\pi$ that quantifies the usefulness of the knowledge in $\pi$ for a particular task.
We focus on the class of risk functions induced by loss functions $l:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$, defined on two variables $x$ and $a$, where $x$ denotes an outcome or state and $a$ denotes an action. Such losses allow us to quantify how useful the knowledge encoded by a partition $\pi$ is by optimizing $l(x,a)$ with respect to $a$ under the knowledge in $\pi$. Throughout this work, we make the common assumptions that $\mathcal{X}$ and $\mathcal{A}$ are standard Borel spaces, $l$ is a jointly measurable function on $\mathcal{X}\times\mathcal{A}$, and all measurability statements are understood modulo $\mathbb{P}$-null sets.
Let us first consider the two extreme cases of full knowledge about $X$, encoded by the partition $\pi(X)=\{X^{-1}(x)\}_{x}$, and no knowledge about $X$, encoded by the partition $\bot\coloneqq\{\Omega\}$. Since knowledge of the realization $x$ of $X$ enables pointwise optimization of the loss $a\mapsto l(x,a)$, whereas without any knowledge only the expected loss $a\mapsto\mathbb{E}[l(X,a)]$ can be optimized, we have
$\mathbb{E}\Big[\inf_{a\in\mathcal{A}}l(X,a)\Big]\leq\inf_{a\in\mathcal{A}}\mathbb{E}[l(X,a)].$ (6)
Some knowledge about $X$, realized by an arbitrary partition $\pi$ of $\Omega$, can lead to an expected loss between these two extremes. The key observation is that a decision-maker with state of knowledge $\pi$ can only choose actions that are compatible with the distinctions available under $\pi$, i.e., actions that are $\sigma(\pi)$-measurable. The two extremes in (6) correspond to the coarsest and finest possible knowledge: under no knowledge only constant actions ($\sigma(\bot)$-measurable actions) are possible, whereas under full knowledge, the action can be a function of $X$, i.e., it is $\sigma(X)$-measurable. This motivates the following definition for arbitrary partitions $\pi$ of $\Omega$.
Definition 3 (Risk).
Given a random variable $X:\Omega\to\mathcal{X}$ and a measurable partition $\pi$ of $\Omega$, let the risk of $\pi$ be defined by
$R(\pi)=\inf_{A\ \sigma(\pi)\text{-meas.}}\mathbb{E}[l(X,A)],$ (7)
where the minimization is over $\sigma(\pi)$-measurable random variables $A:\Omega\to\mathcal{A}$. For random variables $Y$, we write $R(Y)$ for the risk $R(\pi(Y))$.
As we show in the proof of Proposition 1 in Appendix -A, for the special case $\pi=\pi(X)$, we obtain
$R(X)=\mathbb{E}\Big[\inf_{a\in\mathcal{A}}l(X,a)\Big],$
from which follows that for any partition $\pi$ of $\Omega$ we have
$R(\bot)\geq R(\pi)\geq R(X).$ (8)
In particular, once a loss function $l$ is given, we can compare knowledge states $\pi$ based on how well their blocks align with the blocks of $\pi(X)$ in terms of optimizing $l(X,\cdot)$.
II-D Information as risk reduction
First, consider the case of total risk reduction, from no knowledge to full knowledge about $X$, i.e., the difference between the two sides of (6), which we identify as the entropy of $X$,
$H(X)\coloneqq\inf_{a\in\mathcal{A}}\mathbb{E}\big[l(X,a)\big]-\mathbb{E}\Big[\inf_{a\in\mathcal{A}}l(X,a)\Big].$ (9)
Hence, $H(X)$ is the average difference in loss between not knowing and knowing $X$, representing the total uncertainty in $X$ with respect to the loss $l$. Different loss functions may capture different aspects of uncertainty. For example, if the loss function is taken to be the square error, then $H$ is the variance, measuring uncertainty in the sense of the average square distance to the expected value, whereas for the log loss we recover Shannon entropy (see Appendix -B). Note that (9) differs from the decision-theoretic notion of entropy in [35] by subtracting the average loss incurred even if $x$ is known perfectly. In many practical examples, $l$ is chosen such that this term vanishes, by ensuring that $l(x,a)\geq 0$ with equality at the minimum. However, in principle, $l$ can have arbitrary minimal values, which do not affect the value of $H(X)$ in (9). Including this term in the definition makes it possible to treat entropy and information simply as two cases of risk reduction between different knowledge states, as we will see in the following.
Since entropy (9) is the difference of the risks $R(\bot)$ and $R(X)$, inequality (8) suggests to decompose $H(X)$ into two non-negative terms,
$H(X)=\underbrace{R(\bot)-R(\pi)}_{\text{Information }I(X;\pi)}+\underbrace{R(\pi)-R(\pi(X)),}_{\text{Conditional entropy }H(X|\pi)}$
where we identify the risk reduction from some knowledge $\pi$ to full knowledge $\pi(X)$ as the conditional entropy of $X$ given $\pi$, and the risk reduction from no knowledge $\bot$ to some knowledge $\pi$ as information in $\pi$ about $X$. In particular, the usual decomposition of entropy into mutual information and conditional Shannon entropy, $H(X)=I(X;Y)+H(X|Y)$, is a basic feature of our general notion of information for arbitrary loss functions, which we specify in the following.
Figure 3: Schematic illustration of how specifying a loss projects the knowledge lattice on $\Omega$ to the real line of risks, giving rise to informational quantities, such as entropy $H(X)$, conditional entropy $H(X|\pi)$ and information $I(X;\pi)$, through risk reduction from one knowledge state to another.
Definition 4 (Uncertainty, entropy, and information).
Let $X:\Omega\to\mathcal{X}$ be a random variable and let $\pi$ and $\pi^{\prime}$ be two not necessarily comparable partitions of $\Omega$. The uncertainty reduction about $X$ from $\pi$ to $\pi^{\prime}$ (with respect to $l$) is
$U_{\pi\to\pi^{\prime}}(X)\coloneqq R(\pi)-R(\pi^{\prime})$
$\quad\ \ =\inf_{A\,\sigma(\pi)\text{-meas.}}\mathbb{E}[l(X,A)]-\inf_{A\,\sigma(\pi^{\prime})\text{-meas.}}\mathbb{E}[l(X,A)]\,.$ (10)
Furthermore, we distinguish the following quantities.
  1. $(i)$
    The entropy of $X$ conditioned on $\pi$ is given by
    $H(X|\pi)\coloneqq U_{\pi\to\pi(X)}(X)\,$, (11)
    i.e., conditional entropy is the risk reduction from some knowledge state $\pi$ to full knowledge $\pi(X)$. The case $\pi{=}\bot$ corresponds to the total risk reducible by observing $X$, which is the entropy (9) of $X$,
    $H(X)=H(X|\bot)=U_{\bot\to\pi(X)}(X)\,.$ (12)
    Moreover, if $Y$ denotes a random variable with partition $\pi$, then we write $H(X|Y)\coloneqq H(X|\pi(Y))$.
  2. $(ii)$
    The information about $X$ in $\pi^{\prime}$ conditioned on $\pi$ is
    $I(X;\pi^{\prime}|\pi)\coloneqq U_{\pi\to\pi\vee\pi^{\prime}}(X).$ (13)
    For $\pi=\bot$ we obtain the information about $X$ in $\pi^{\prime}$,
    $I(X;\pi)\coloneqq I(X;\pi|\bot)=U_{\bot\to\pi}(X),$ (14)
    since $\bot\vee\pi=\pi$. We write $I(X;Y)\coloneqq U_{\bot\to\pi(Y)}(X)$ and, moreover, since $\pi(Y,Z)=\pi(Y)\vee\pi(Z)$, we also write
    $I(X;Y|Z)\coloneqq U_{\pi(Z)\to\pi(Y,Z)}(X)$
    for random variables $Y,Z$.
Without further assumptions on $\pi$ and $\pi^{\prime}$, we cannot say whether $U_{\pi\to\pi^{\prime}}$ is positive or negative, corresponding to a decrease or increase in uncertainty, respectively. The special cases $(i)$ and $(ii)$, however, are non-negative, as we show in Proposition 1 below. Concrete examples for common loss functions, including square error, log loss, general Bregman losses, and proper scoring rules, are discussed in Appendix -B.
Note that $I(X;Y)$ is inherently asymmetric with respect to interchanging $X$ and $Y$, since $X$ directly determines the loss $l(X,\cdot)$, whereas $Y$ generally contributes only through the partition $\pi(Y)$. This asymmetry reflects the fact that $I(X;Y)$ quantifies the knowledge gain about $X$ when moving from no knowledge to observing $Y$, whereas $I(Y;X)$ quantifies the knowledge gain about $Y$ when moving from no knowledge to observing $X$. These two quantities are therefore conceptually distinct and coincide only for specific choices of the objective $l$ with respect to which the knowledge gain is measured, such as the log loss.
The following properties of entropy and information are shown in Appendix -A.
Proposition 1.
For any random variable $X:\Omega\to\mathcal{X}$, we have:
  1. $(i)$
    Maximality of $\pi(X)$: $U_{\pi\to\pi(X)}(X)\geq 0$ for any $\pi$, in particular $H(X|\pi)$ is non-negative.
  2. $(ii)$
    Non-negativity: If $\pi\preceq\pi^{\prime}$, then $R(\pi^{\prime})\leq R(\pi)$ (i.e., R-R is a monotone w.r.t. $\preceq$), and therefore $U_{\pi\to\pi^{\prime}}\geq 0$. In particular, $I(X;\pi^{\prime}|\pi)$ is non-negative.
  3. $(iii)$
    No uncertainty: If $\mathbb{P}_{X}=\delta_{x}$ (the Dirac measure at $x$) for some $x\in\mathcal{X}$ then $H(X)=0$.
  4. $(iv)$
    Independence: If $X$ and $Y$ are stochastically independent then
    $I(X;Y)=0\,,\quad H(X|Y)=H(X)\,.$
  5. $(v)$
    Chain rule: $I(X;(Y,Z))=I(X;Y)+I(X;Z|Y)$, and in general, for any sequence $(\pi_{n})_{n=0}^{N}$ with $n\mapsto R(\pi_{n})$ being non-increasing,
    $U_{\pi_{0}\to\pi_{N}}(X)=\sum_{n=1}^{N}U_{\pi_{n-1}\to\pi_{n}}(X)$ (15)
    is a non-negative decomposition.
Proposition 1 shows that many well-known properties of Shannon entropy and information are retained when allowing the loss function to be arbitrary. Not only that, but in Section III, we show that the underlying structure of the knowledge lattice enables a principled treatment of partial information decompositions (PIDs) using (15).
III Partial information decomposition
According to the chain rule in Proposition 1, information measures decompose into multiple non‑negative terms when evaluated along chains of states with increasing knowledge. A prominent example of such a decomposition for Shannon information (i.e., log loss) is partial information decomposition (PID) [83], which has attracted considerable attention in the literature.
The practical relevance of PIDs lies in the ability to distinguish qualitatively different modes of multivariate dependence. While Shannon information aggregates all dependence into a single scalar, PIDs separate redundant contributions (shared among the sources), unique contributions (source-specific), and synergistic contributions (obtainable only from joint observations). This distinction is crucial in fields such as neuroscience [82, 18], computational biology [74, 84], machine learning [22], and complex systems analysis [50, 57], where understanding the shared, unique, and synergistic contributions of signals can change both interpretation and modeling strategy. For example, in neural coding, failing to separate synergistic from redundant contributions can lead to fundamentally different conclusions about whether a neural population encodes information collectively or individually [70, 48, 74]. Due to this sustained interest, after the seminal paper [83] there have been many contributions to the foundations of PID, [33, 7, 43, 66, 3, 52], to name a few.
However, recently it has been shown that most current PID formulations are incompatible with core information-theoretic properties such as the chain rule and non-negativity [54]. In the following, we offer a possible solution to this issue. We show how PID‑type decompositions arise naturally within our framework by interpreting redundant, unique, and synergistic information in terms of the structure of the underlying knowledge lattice, thereby providing links to existing concepts from probability theory, such as Aumann’s common knowledge. In particular, we challenge one basic assumption of PIDs, namely that unique contributions are independent of already available knowledge (see the discussion in the following section). Moreover, while most PID formulations are specific to Shannon information, analogous decomposition questions exist for other uncertainty measures such as variance, where so-called variance decompositions [49] face similar conceptual difficulties. Our approach yields partial information decompositions for any loss function, thereby unifying these settings.
III-A Current PID assumptions
First, consider the case of three random variables $X$, $X_{1}$, and $X_{2}$, where $X$ is considered to be the quantity of interest, and $X_{i}$ are examined with respect to their information about $X$. Using the chain rule of mutual information, we can decompose total Shannon information $I_{S}(X;(X_{1},X_{2}))$ between $X$ and $(X_{1},X_{2})$ into the Shannon information in $X_{1}$ about $X$, $I_{S}(X;X_{1})$, plus the Shannon information in $X_{2}$ about $X$ given $X_{1}$, $I_{S}(X;X_{2}|X_{1})$. PIDs attempt to further decompose these terms into redundant, unique, and synergistic information,
$I_{S}(X;(X_{1},X_{2}))=R_{1,2}+U_{1}+U_{2}+S_{1,2},$ (16)
where the redundant and synergistic terms, $R_{1,2}$ and $S_{1,2}$, are symmetric with respect to interchanging $X_{1}$ and $X_{2}$, whereas the unique terms are not, and all terms are assumed to be non-negative. Note that $U_{1}$, the unique information of $X_{1}$ about $X$, also depends on the random variable $X_{2}$ with respect to which it has to be unique, and vice versa. The decomposition (16) results from the following assumptions on the single-variable information terms:
$I_{S}(X;X_{i})$ $=R_{1,2}+U_{i}$ (17)
$I_{S}(X;X_{i}|X_{j})$ $=U_{i}+S_{1,2}\,$, (18)
where $i,j\in\{1,2\}$ with $i\neq j$. Notably, the same unique information $U_{i}$ appears in both quantities, no matter whether the corresponding variable is observed without any other information or whether there is already information about $X$ available. In fact, as we argue below, this path independence is only appropriate at the level of information content (i.e., knowledge states), not at the level of measuring information. Despite this constraint, there are still too many degrees of freedom left even if both (17) and (18) are satisfied, because the four equations are not linearly independent and therefore do not permit determining the four terms $R_{1,2},U_{1},U_{2},S_{1,2}$.
Therefore, there are various proposals in the literature that each impose a different additional assumption in order to specify how either redundant, unique, or synergistic information should be identified, so-called PID-inducing conditions [36], from which the full decomposition then follows through (17) and (18). For example, Williams and Beer [83] define redundant information as the average of the minimum information about outcomes $X=x$ provided by each of the variables $X_{i}$. Griffith and Koch [33] quantify synergistic information, motivated by [55], as the difference between the total Shannon information $I_{S}(X;(X_{1},X_{2}))$ when evaluated using the actual joint $\mathbb{P}_{X,X_{1},X_{2}}$ and when evaluated using the joint $q_{X,X_{1},X_{2}}$ that minimizes the total information under the constraint that the marginals $q_{X,X_{i}}$ are the same as $\mathbb{P}_{X,X_{i}}$. Similarly, Bertschinger et al. [7] define unique information $U_{i}$ as the minimum of $I_{S}(X;X_{i}|X_{j})$ for $j\neq i$ over all joint distributions preserving the marginals $\mathbb{P}_{X,X_{i}}$. Note that, for ease of presentation, we have discussed the case of two sources here, but PIDs are often extended to the case of arbitrarily many sources $X_{1},\dots,X_{n}$, with increasingly many terms in the decomposition, e.g., 18 terms for $n=3$ [33] (cf. Appendix -E, where we show one path through the knowledge lattice resulting in a finer PID of 32 terms).
Figure 4: Partial information decomposition for three variables. When traversing the knowledge lattice along the shown path, the total information $I(X;(Y,Z))$ decomposes into four terms, corresponding to redundant, unique, conditional unique, and synergistic information.
A central limitation in much of the current PID literature is that informational quantities themselves are treated as primitive objects that can be combined through addition. However, in contrast to the knowledge states $\pi$, information is fundamentally a scalar quantity, whose additivity is generally justified only through the chain rule. As a consequence, the intuitive notion of one information term being “contained” in another has limited formal applicability. For example, this issue applies to how current PID formulations consider unique information to be the same no matter which knowledge is already available, i.e., whether the unique term is part of $I_{S}(X;X_{i})$ or $I_{S}(X;X_{i}|X_{j})$ (see (17) and (18) above). Similarly, in a recent approach [69], redundant information is defined via a collection of information terms $I_{i}$ whose sum is required to be “contained” in each single-variable information term $I_{S}(X;X_{i})$. Yet, there is no technical justification for assuming an additivity property of single-variable information terms beyond the identities imposed by the chain rule. The same applies to the work in [37, 36], where so-called part-whole relationships are considered the fundamental principle of how PIDs are constructed in general, which inherently treat informational terms as quantities that are “part” of other quantities.
The most prominent example of this type of treatment is the extensive use of Venn diagrams in the current PID literature, known as PI-diagrams [33]. The interpretation of standard Venn diagrams representing relationships between informational quantities through overlapping two-dimensional regions is conceptually difficult, both because they may contain regions corresponding to negative terms and because they suggest inappropriate set-theoretic analogies [51]. While PI-diagrams are defined to contain only non-negative regions, they still give a false set-theoretic intuition for quantities that are purely algebraic. The same applies to variance decompositions [49], which also use Venn diagrams to represent overlap between contributions. Informational quantities should instead be understood as outcomes of transitions between states of knowledge, rather than as objects that themselves admit a compositional structure. It is the states of knowledge $\pi$ that constitute the primitive structures that can be combined through meets and joins, thereby capturing the notions of redundancy, synergy, and unique information.
III-B Redundancy, uniqueness, and synergy
Recall that the risk $R(Y)$ is the minimal expected loss $\mathbb{E}[l(X,A)]$ achievable by actions $A$ measurable with respect to $\sigma(Y)$. Since the expected loss depends on $Y$ only through the expectation of $\mathbb{E}[l(X,A)|Y]$ due to the tower property of conditional expectations, $R(Y)$ only depends on the random conditional probability measure $\mathbb{P}_{X\mid Y}$ and not on the full observation $Y$ itself. Equivalently, the infimum over $\sigma(Y)$‑measurable actions may be restricted to the smaller $\sigma$‑algebra $\sigma(\mathbb{P}_{X\mid Y})$. In terms of partitions, this corresponds to replacing the observation partition $\pi(Y)$ by the coarser partition
$\pi_{Y}\coloneqq\pi(\mathbb{P}_{X\mid Y})\preceq\pi(Y),$
which identifies outcomes of $Y$ that induce the same conditional law of $X$. We summarize this in the following proposition.
Proposition 2.
For any random variable $Y$, we have
$R(Y)=R(\pi_{Y}).$
Note that, if $\mathbb{P}_{X}$ is viewed as the (constant) random variable $\omega\mapsto\mathbb{P}_{X}$, then $\pi(\mathbb{P}_{X})=\bot$, and moreover, for $Y=X$, we have $\pi_{X}=\pi(\delta_{X})=\pi(X)$, where $\delta_{x}$ denotes the Dirac measure centered at $x\in\mathcal{X}$.
Figure 5: Knowledge lattices for the three most basic examples of three variables $X,Y$ and $Z$, each example focussing on a different type of information: $\mathsf{UNQ}$ only contains unique information, $\mathsf{RDN}$ only redundant information, and $\mathsf{XOR}$ only contains synergistic information. The nodes, each representing a partition of $\Omega=\{(0,0),(0,1),(1,0),(1,1)\}$, are visualized using different textures for different blocks of the corresponding partition. Next to the partitions we display the associated distributions over $X$. The non-zero risk reduction terms are calculated for the log loss, mean squared error and zero-one loss, discussed in Appendix -B.
By Proposition 2, we may replace $\pi(Y)$ by $\pi_{Y}$ in the definition of entropy and information, i.e.
$H(X|Y)$ $=R(\pi_{Y})-R(X),$
$I(X;Y)$ $=R(\bot)-R(\pi_{Y}),$
$I(X;Z|Y)$ $=R(\pi_{Y})-R(\pi_{Y,Z})\,.$
We can therefore interpret a path of risk reduction from no knowledge to partial knowledge to full knowledge about $X$ as a sequence of uncertainty reductions from $\mathbb{P}_{X}$ to $\mathbb{P}_{X\mid Y}$ and finally to $\delta_{X}$ (see also Example 5 in Appendix -B). As we show in the following, when passing to the corresponding knowledge lattice, such a path can be refined further via the meet and join of the associated partitions. This refinement results in a partial information decomposition grounded directly in the structure of the knowledge lattice.
In contrast to partitions induced by observations, for which $\pi(Y,Z)=\pi(Y)\vee\pi(Z)$, the analogous identity for $\pi_{Y,Z}$ generally fails, i.e.,
$\pi(\mathbb{P}_{X|Y,Z})\neq\pi(\mathbb{P}_{X|Y})\vee\pi(\mathbb{P}_{X|Z})\,.$
The difference of the corresponding risks, in fact, is what we call synergistic information (see Definition 5 below), because the term synergy describes exactly what is gained when the observations $Y$ and $Z$ are available together in comparison to having observed $Y$ and $Z$ separately, which is in alignment with most of the PID literature.
The other two types of information that are part of a PID also have natural candidates on the knowledge lattice. Redundant information is simply the risk reduction from no knowledge to common knowledge [2] of $\pi_{Y}$ and $\pi_{Z}$, represented by $\pi_{Y}\wedge\pi_{Z}$. Unique information in $Y$ about $X$ is the risk reduction associated with $\pi_{Y}$, where we additionally distinguish whether the unique term is part of $I(X;Y)$, where it is risk reduction from common knowledge to $\pi_{Y}$, or whether it is part of $I(X;Y|Z)$, where it is risk reduction from $\pi_{Z}$ to $\pi_{Y}\vee\pi_{Z}$ (conditional unique information).
Definition 5 (Redundancy, uniqueness, and synergy).
Let $(\Omega,\Sigma,\mathbb{P})$ be a probability space, let $l:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$ be a measurable function, and let $X:\Omega\to\mathcal{X}$ be a random variable. In the following, $Y$ and $Z$ denote two random variables that are considered with respect to what their observations tell about $X$.
  1. $(i)$
    The redundant information in $Y$ and $Z$ about $X$ is given by the risk reduction from no knowledge to common knowledge,
    $I_{r}(X;(Y,Z))\coloneqq R(\bot)-R(\pi_{Y}\wedge\pi_{Z})\,.$ (19)
  2. $(ii)$
    The unique information in $Y$ about $X$ (without $Z$) is the risk reduction from common knowledge $\pi_{Y}\wedge\pi_{Z}$ to $\pi_{Y}$ (or, equivalently $\pi(Y)$),
    $I_{u}(X;Y{\setminus}Z)\coloneqq R(\pi_{Y}\wedge\pi_{Z})-R(\pi_{Y})\,$, (20)
    where $\setminus$” indicates that this is the unique information in $Y$ without the common knowledge that is also part of $Z$.
  3. $(iii)$
    The conditional unique information in $Y$ about $X$ given $Z$ is the risk reduction from $\pi_{Z}$ (or, equivalently $\pi(Z)$) to the join of $\pi_{Y}$ and $\pi_{Z}$,
    $I_{cu}(X;Y|Z)\coloneqq R(\pi_{Z})-R(\pi_{Y}\vee\pi_{Z})\,$, (21)
    where “||” indicates that this is the unique information in $Y$ given that $Z$ is already known.
  4. $(iv)$
    The synergistic information of $Y$ and $Z$ about $X$ is the risk reduction from the join of $\pi_{Y}$ and $\pi_{Z}$ to $\pi_{Y,Z}$,
    $I_{s}(X;(Y,Z))\coloneqq R(\pi_{Y}\vee\pi_{Z})-R(\pi_{Y,Z})\,.$ (22)
Figure 6: Knowledge lattices for combinations of the basic examples in Figure 5. Each has a PID (25) with two non-zero terms: $\mathsf{RDN}\otimes\mathsf{UNQ}$ has redundant and unique contributions, $\mathsf{UNQ}\otimes\mathsf{XOR}$ has unique and synergistic information, and $\mathsf{RDN}\otimes\mathsf{XOR}$ has redundant and synergistic terms.
Note that conditional unique information is not part of existing PID formulations in the literature. It appears in our decomposition (Proposition 3 below), because transitioning from common knowledge to $\pi_{Y}$ is generally not equivalent to transitioning from $\pi_{Z}$ to $\pi_{Y}\vee\pi_{Z}$. Consequently, unique information has a different value depending on which information is already available. However, all existing PID approaches based on the foundational assumptions in [83] use the same unique information term in both situations. This discrepancy reflects the more general issue discussed earlier: informational quantities are not inherently stable objects themselves but are the results of transitions between the underlying states of knowledge. In particular, the resulting PID depends on the path taken through the knowledge lattice.
The non-negativity of the redundant, unique, and conditional unique information measures defined in Definition 5 follows directly from the corresponding partitions being comparable with respect to refinement. That synergistic information is also non-negative is shown in Appendix -D as part of the proof of the following proposition.
Proposition 3 (Partial information decomposition).
The quantities defined in Definition 5 are all non-negative and we have
$I(X;Y)$ $=I_{r}(X;(Y,Z))+I_{u}(X;Y{\setminus}Z)$ (23)
$I(X;Z|Y)$ $=I_{cu}(X;Z|Y)+I_{s}(X;(Y,Z))\,.$ (24)
In particular, we obtain
$I(X;(Y,Z))\ =$ $\ I_{r}(X;(Y,Z))+I_{u}(X;Y{\setminus}Z)$
$+I_{cu}(X;Z|Y)+I_{s}(X;(Y,Z))\,$, (25)
i.e., a partial information decomposition for the three variables $X$, $Y$ and $Z$ based on the knowledge lattice on $\Omega$.
The PID in (25) corresponds to first observing $Y$ and then $Z$, i.e., the path through the lattice shown in Figure 4. As can be seen from the examples in the following section, the terms in these decompositions are generally non-trivial.
Most of the PID literature also considers decompositions of $I(X;(X_{1},\dots,X_{n}))$ for $n>2$ [83, 33]. To this end, a simplified set-based notation of the form $\{i\cdots j,\dots,k\cdots l\}$ is commonly used. Here, commas indicate redundant contributions, joint indices without commas denote synergistic terms, whereas unique information terms do not include the other variables. For example, for $n=2$, the four PID components in (16) are denoted by $\{1,2\}$ for the redundant information, $\{1\}$ and $\{2\}$ for the respective unique information terms, and $\{12\}$ for the synergistic information. If we write $\{2|1\}$ for conditional unique information (21) of $X_{2}$ about $X$ given $X_{1}$ then our proposed decomposition (25) consists of the four terms $\{1,2\},\{1\},\{2|1\},\{12\}$.
For $n=3$, the PID literature considers 18 distinct informational terms [33]. Within the knowledge lattice, many paths exist from $\bot$ to $\pi_{X_{1},X_{2},X_{3}}$ that each induce a decomposition of $I(X;(X_{1},X_{2},X_{3}))$, also with more than 18 terms. Figure 9 in the appendix illustrates one such path consisting of 32 non-trivial contributions.
III-C Examples
In this section, we provide basic examples motivated by the examples in [33], including cases that have only synergistic $(a)$, unique $(b)$, and redundant $(c)$ information, examples that combine two of these $(d)$, and a final example showing the necessity of path-dependence $(e)$.
$(a)$ The most obvious example is $X=\mathsf{XOR}(Y,Z)$, which requires full knowledge about both $Y$ and $Z$ in order to know anything about $X$, and so there is only synergistic information. As can be seen in Figure 5 on the right, this is realized by all partitions $\pi_{Y}\wedge\pi_{Z}$, $\pi_{Y}$, $\pi_{Z}$, and $\pi_{Y}\vee\pi_{Z}$ being trivial, so that only the final step to $\pi_{Y,Z}=\pi(\mathbb{P}_{X|Y,Z})$ results in non-zero information $I_{s}(X;(Y,Z))=R(\bot)-R(\pi_{Y,Z})$.
$(b)$ Similarly, the example $X=\mathsf{UNQ}(Y,Z)$ requires both $Y$ and $Z$ to determine $X$ fully, but knowing either of the two already reduces the possible values of $X$ to a smaller range, e.g., $X\in\{2,3\}$ for $Y=1$. Figure 5 also shows the corresponding distributions $\mathbb{P}_{X|Y}$ and $\mathbb{P}_{X|Z}$. It follows that $\pi_{Y}\vee\pi_{Z}$ is already enough to fully specify $X$ so that the synergistic term is zero, i.e., it is sufficient to know the distributions rather than the actual outcomes, which, for example, is not the case for $\mathsf{XOR}$. Moreover, there is no redundant information, because there is no common knowledge between $\pi_{Y}$ and $\pi_{Z}$.
$(c)$ In contrast, the example $X=\mathsf{RDN}(Y,Z)$ only requires knowing one of $Y$ or $Z$ to determine $X$, so that the information is fully redundant because $\pi_{Y}\wedge\pi_{Z}=\pi_{Y,Z}$.
$(d)$ In Figure 6, we illustrate the possible combinations of two of these basic examples, which then only have one trivial type of information instead of two. For example, $\mathsf{UNQ}\otimes\mathsf{XOR}$ consists of four copies of $\mathsf{XOR}$, where getting to know $Y$ or $Z$ narrows down the possibilities to two copies of $\mathsf{XOR}$ and the joint knowledge $\pi_{Y}\vee\pi_{Z}$ determines the final copy of $\mathsf{XOR}$, while the joint observation $(Y,Z)$ is required to determine the final value of $X$.
Figure 7: An example showing the path dependence of unique information terms (cf. $(e)$ in Section III-C). $\mathsf{UNQ_{3}}$ is a modified version of $\mathsf{UNQ}$ from Figure 5 by removing the case $(Y,Z)=(1,1)$. Similar to $\mathsf{UNQ}$, there are no redundant and synergistic contributions, and moreover, $I_{u}(X;Y{\setminus}Z)\neq I_{cu}(X;Y|Z)$.
$(e)$ Consider the variation of $\mathsf{UNQ}$ shown in Figure 7, termed $\mathsf{UNQ}_{3}$, which removes the case $(Y,Z)=(1,1)$ from $\mathsf{UNQ}$, so that $\Omega$ is now a three-element set. As can be seen from the distributions and corresponding partitions, the two paths from common knowledge $\pi_{Y}\wedge\pi_{Z}$ to joint knowledge $\pi_{Y}\vee\pi_{Z}$ via $\pi_{Y}$ and $\pi_{Z}$ are symmetric, so that $I_{u}(X;Y{\setminus}Z)=I_{u}(X;Z{\setminus}Y)$ and $I_{cu}(X;Y|Z)=I_{cu}(X;Z|Y)$. However, the unique and conditional unique information terms are not the same,
$I_{u}(X;Y{\setminus}Z)\neq I_{cu}(X;Y|Z).$
The same is true when switching the roles of $Y$ and $Z$, as explicitly calculated for Shannon information, mean square error, and the zero-one loss in Figure 7. In particular, regardless of our precise knowledge-lattice-based formulation, if we agree that there are no redundant and synergistic contributions, then it already follows from the decomposition (17)-(18) of $I(X;X_{1})$ and $I(X;X_{1}|X_{2})$ that the two appearances of $U_{1}$ cannot be the same.
The last example has shown that unique information is path dependent, i.e., that it depends on the already available knowledge: while the knowledge content of $Y$, represented by $\pi_{Y}$, is independent of getting to know $Z$ before or after observing $Y$, the order does make a difference when measuring this content through a specified loss function.
III-D Finer decompositions
Since our partial information decomposition is purely based on the knowledge lattice and the chain rule, we can create arbitrary decompositions given suitable partitions. One such refinement of the usual decomposition of $I(X;(Y,Z))$ into the four terms discussed above would be to split up redundant information and conditional unique information even further, by including the meet with $\pi_{Y,Z}$ as a separate step which results in a decomposition that allows to further disambiguate informational contributions to the total information.
In particular, instead of moving from $\bot$ directly to $\pi_{Y}\wedge\pi_{Z}$, we could include the meet with $\pi_{Y,Z}$ as a middle step. More precisely,
$I_{dr}(X;(Y,Z))$ $\coloneqq R(\bot)-R(\pi_{Y}\wedge\pi_{Z}\wedge\pi_{Y,Z}),$
$I_{pr}(X;(Y,Z))$ $\coloneqq R(\pi_{Y}\wedge\pi_{Z}\wedge\pi_{Y,Z})-R(\pi_{Y}\wedge\pi_{Z}),$
where $I_{dr}$ might for example be called deep redundancy, capturing the knowledge common in $Y$, $Z$, and $(Y,Z)$, and $I_{pr}$ describes pairwise redundancy, capturing the part that is common in $Y$ and $Z$ but not in the joint observation $(Y,Z)$. Then we obtain
$I_{r}(X;(Y,Z))=I_{dr}(X;(Y,Z))+I_{pr}(X;(Y,Z)),$ (26)
and both terms are non-negative due to the chain rule.
Similarly, for conditional unique information, instead of moving from, say, $\pi_{Y}$, directly to $\pi_{Y}\vee\pi_{Z}$, we could again include the meet with $\pi_{Y,Z}$ as a middle step. We write
$I_{acu}(X;Z|Y))$ $\coloneqq R(\pi_{Y})-R(\pi_{Y}\vee(\pi_{Z}\wedge\pi_{Y,Z})),$
$I_{rcu}(X;Z|Y))$ $\coloneqq R(\pi_{Y}\vee(\pi_{Z}\wedge\pi_{Y,Z}))-R(\pi_{Y}\vee\pi_{Z}),$
for aligned conditional unique information $I_{acu}$, capturing the extra gain when adding the common knowledge between $\pi_{Z}$ and $\pi_{Y,Z}$, and residual conditional unique information $I_{rcu}$, capturing the extra knowledge in $\pi_{Z}$ that is not in $\pi_{Y,Z}$. Then
$I_{cu}(X;Z|Y))=I_{acu}(X;Z|Y))+I_{rcu}(X;Z|Y)\,.$ (27)
There are examples that only contain deep redundancy or only pairwise redundancy. For instance, our previous redundancy examples $\mathsf{RDN}$, $\mathsf{RDN}\otimes\mathsf{UNQ}$, and $\mathsf{RDN}\otimes\mathsf{XOR}$ all only have deep redundancy, whereas one obtains only pairwise redundancy by modifying the latter so that the second $\mathsf{XOR}$ has the values $(0,2,2,0)$ instead of $(2,3,3,2)$.
For the refinement of conditional unique information, we show in Figure 8 the knowledge lattice of $\mathsf{AND}$, whose PID terms are debated in the literature [33, 3]. Our refined PID offers a possible answer: Similarly to $\mathsf{UNQ}$, there is no redundant or synergistic information. However, in contrast to $\mathsf{UNQ}$, aligned conditional unique information is zero, because there is no common knowledge shared between, say, $\pi_{Z}$ and $\pi_{Y,Z}$, and so all that is gained when moving from $\pi_{Y}$ to $\pi_{Y}\vee\pi_{Z}$ is residual conditional unique information.
In particular, these examples show that the extra terms in the refined decompositions (26) and (27) are generally non-trivial. Moreover, these are just examples of possible refinements of PIDs. One could also move from $\pi_{Y}\wedge\pi_{Z}$ first to $\pi_{Y}\wedge(\pi_{Z}\vee\pi_{Y,Z})$ and then to $\pi_{Y}$, just to name another possible extension. These types of refinements also apply to the case of three sources, for which we have shown one particular risk sequence in Figure 9 in the appendix.
Figure 8: Knowledge lattices for $\mathsf{AND}$ and $\mathsf{UNQ}$ under the finer decomposition of conditional unique information into aligned and residual conditional unique information. For $\mathsf{AND}$, we have $I_{acu}=0$ and $I_{rcu}\neq 0$, whereas for $\mathsf{UNQ}$, we have $I_{acu}\neq 0$ and $I_{rcu}=0$.
IV Conclusion
The concept of information plays a central role across science and engineering, yet it is often implicitly identified with Shannon entropy or mutual information. While these quantities originate from optimal coding with respect to message length, their broader use as generic uncertainty measures can obscure the distinction between informational content, understood as what is known, and information as a scalar quantity, understood as how knowledge is evaluated. In this work, we made this distinction explicit and provided a principled separation between informational content and its numerical assessment. In particular, our goal is not to argue against the use of Shannon information, but rather to expose the general knowledge-lattice structure that underlies informational quantities, from which, for example, principled decompositions of information follow naturally.
More precisely, we identified a general structural description of information and entropy in terms of transitions between knowledge states and their quantification through loss-based optimization. Knowledge is modeled through partitions of the underlying probability space, or equivalently by the corresponding $\sigma$-algebras of knowable sets, forming a knowledge lattice ordered by refinement. Scalar informational quantities are obtained once a loss function is specified, which enables us to measure the value of knowledge states via their achievable risks. Accordingly, entropy and information are unified as instances of risk reduction relative to different states of knowledge.
While many well-known information and uncertainty measures fit into the risk-based structure adopted here, such as those induced by Bregman losses (including Shannon information and variance) and proper scoring rules (e.g., Tsallis and Brier scores), it does not include all generalized information measures. For example, Fisher information quantifies parametric sensitivity rather than uncertainty reduction between knowledge states. Moreover, it remains unclear whether Rényi entropy admits a representation as expected loss minimization. One possible generalization would be to replace risk differences by arbitrary divergences between probability distributions, such as $f$-divergences [19], which would still allow developing informational quantities based on the knowledge lattice, but may come at the cost of losing certain structural properties such as the chain rule.
In continuous settings, the possibility of unbounded uncertainty reduction is due the use of pointwise loss functions defined on continuous action spaces. This is in alignment with the divergence of Shannon entropy under increasingly fine discretizations. However, since information is defined directly in terms of risk reduction rather than as a difference of entropies, we avoid the need for cancellations of infinities in order to obtain finite quantities.
A key application of our approach is the decomposition of informational quantities. By grounding PID in the structure of the knowledge lattice, we arrive at straightforward definitions of redundancy, unique information, and synergy based on meets and joins of knowledge states. While our definition of synergy is closely related to the PID literature, capturing the difference between marginals and full conditionals, redundancy is often less obvious to identify. On the knowledge lattice, it simply corresponds to Aumann’s established notion of common knowledge. The unique contributions are necessarily path-dependent, taking different values depending on the knowledge already available. This path dependence is not a deficiency, but a consequence of the fact that information is an outcome of transitions between knowledge states rather than an object with an intrinsic compositional structure.
Several lattice-based approaches in the PID literature, most prominently the redundancy lattice [83] and more recent parthood lattices [37], organize decomposition terms under the assumption that informational quantities themselves admit part-whole relationships. These constructions are therefore combinatorial rather than epistemic in nature, serving primarily to identify and index terms rather than to capture relationships between knowledge states. A notable exception is the constraint lattice of [3], whose elements represent probabilistic structures rather than scalar quantities, making it potentially closer in spirit to the knowledge lattice introduced here, albeit without distinguishing redundant and unique information. Investigating formal connections between these frameworks is an interesting direction for future work. More broadly, our approach suggests that the foundations of information decomposition should be sought in the structure of probability spaces themselves, in the meets, joins, and further refinements of knowledge states, rather than in algebraic relationships among the scalar quantities they induce.
-A Proof of Proposition 1
$(i)$ First of all, we show that
$R(X)=\mathbb{E}\Big[\inf_{a\in\mathcal{A}}l(X,a)\Big].$ (28)
Since $l$ is jointly measurable and $\mathcal{A}$ and $\mathcal{X}$ are standard Borel, by the Jankov-von Neumann theorem [8], the map $x\mapsto\inf_{a\in\mathcal{A}}l(x,a)$ is universally measurable and for every $\varepsilon>0$ there exists a universally measurable selection $f_{\varepsilon}:\mathcal{X}\to\mathcal{A}$ with $l(x,f_{\varepsilon}(x))\leq\inf_{a}l(x,a)+\varepsilon$. In particular, $f_{\varepsilon}(X)$ is $\sigma(X)$-measurable up to $\mathbb{P}$-null sets, and it follows that
$\inf_{A\,\sigma(X)\text{-meas.}}\mathbb{E}[l(X,A)]$ $\leq\mathbb{E}[l(X,f_{\varepsilon}(X))]$
$\leq\mathbb{E}[\inf_{a}l(X,a)]+\varepsilon.$
Since $\varepsilon>0$ was arbitrary, and the opposite inequality is trivial because $\mathbb{E}[l(X,A)]\geq\mathbb{E}[\inf_{a\in\mathcal{A}}l(X,a)]$ for any $A$, this proves (28). Finally, we have
[…]
$(ii)$ If $\pi\preceq\pi^{\prime}$, then $\sigma(\pi)\subseteq\sigma(\pi^{\prime})$, and so if $A$ is $\sigma(\pi)$-measurable, it is also $\sigma(\pi^{\prime})$-measurable, i.e., $R(\pi^{\prime})\leq R(\pi)$ and thus $U_{\pi\to\pi^{\prime}}(X)\geq 0$.
$(iii)$ If $\mathbb{P}_{X}=\delta_{x}$ for some $x\in\mathcal{X}$, then $\mathbb{E}[l(X,a)]=l(x,a)$ and therefore $R(\bot)=R(X)$, i.e., $H(X)=0$.
$(iv)$ If $X$ and $Y$ are independent and $A$ is $\sigma(Y)$-measurable, then, by the Doob-Dynkin Lemma,
$\mathbb{E}[l(X,A)]=\mathbb{E}_{\mathbb{P}_{Y}}[L(\mathbb{P}_{X},f(Y))]$
for some measurable function $f$ and $L(p,a)\coloneqq\mathbb{E}_{p}[l(X,a)]$. It follows that
$\inf_{A\,\sigma(Y)\text{-meas.}}\mathbb{E}[l(X,A)]=\inf_{a}L(\mathbb{P}_{X},a)=\inf_{a\in\mathcal{A}}\mathbb{E}[l(X,a)]$
and therefore $I(X;Y)=0$.
$(v)$ By definition, $I(X;(Y,Z))=U_{\bot\to\pi(Y,Z))}$ and since $\pi_{0}=\bot$, $\pi_{1}=\pi(Y)$, $\pi_{2}=\pi(Y,Z)$ defines a sequence with decreasing risks, $R(\bot)\geq R(Y)\geq R(Y,Z)$, the chain rule for $I(X;(Y,Z))$ follows from the general chain rule (15). The latter is just a telescope sum of terms of the form $R(\pi_{i+1})-R(\pi_{i})$, and each term is non-negative by assumption of $i\mapsto R(\pi_{i})$ being monotonically decreasing.
-B Information for specific losses
In the following, we work out the resulting notions of uncertainty and information for a range of commonly used loss functions. For square error loss, uncertainty reduces to variance and information to explained variance, recovering classical $L^{2}$ decompositions. For log loss, our definitions recover Shannon entropy and mutual information in the discrete case, while in the continuous case entropy becomes infinite but information remains well-defined in terms of expected KL divergences. More generally, for Bregman losses, information coincides with Bregman information (or Jensen gaps), linking our framework to bias-variance decompositions and prediction uncertainty. Finally, for arbitrary losses admitting Bayes acts, information admits an interpretation as the cost of transforming distributions, generalizing the familiar relationship between Shannon information and KL divergence. These examples illustrate that the proposed theory both subsumes classical information measures and extends naturally to objectives beyond coding.
Example 1 (Square error).
The square error on the real line $\mathcal{X}=\mathbb{R}$ is the loss function $l$ defined on $\mathbb{R}^{2}$ (here $\mathcal{A}=\mathcal{X}$) given by
$l(x,a)=(x-a)^{2}.$
The conditional expectation $\mathbb{E}[X|\mathcal{F}]$ is the best $\mathcal{F}$-measurable predictor of $X$ with respect to the square error $l$ [63] (see also (1)). In particular, we obtain
$H(X)$ $=\mathbb{E}[(X-\mathbb{E}[X])^{2}]=\mathbb{V}(X),$
$H(X|Y)$ $=\mathbb{E}[(X-\mathbb{E}[X|Y])^{2}]=\mathbb{E}[\mathbb{V}(X|Y)],$
$I(X;Y)$ $=\mathbb{V}(X)-\mathbb{E}[\mathbb{V}(X|Y)]=\mathbb{V}(\mathbb{E}[X|Y]),$
$I(X;Y|Z)$ $=\mathbb{E}_{Z}[\mathbb{V}_{Y}(\mathbb{E}[X|Y,Z])].$
The variance $\mathbb{V}(X)$ of a random variable $X$ thus corresponds to its total uncertainty measured with respect to the square error. Conditional variance $\mathbb{V}(X|Y)$ is the total uncertainty left when $Y$ is known, whereas $I(X;Y)$ is the variance of the random variable $\mathbb{E}[X|Y]$ resulting from compressing $X$ according to the knowledge in $\sigma(Y)$. Conditional information is a combination of these two when another source of knowledge is taken into account. The analogous results for the square error on $\mathbb{R}^{d}$ follow along the same lines, replacing $(\cdot)^{2}$ by $\|\cdot\|^{2}$.
Similar to how Shannon entropy is the minimal expected codeword length, variance is the optimal square Euclidean distance between $X$ and a single number representing $X$ under no knowledge, capturing the variability of $X$. Conditional variance is an important tool in econometrics, especially in autoregressive conditional heteroskedasticity (ARCH) models [13].
Example 2 (Log loss).
For the discrete case, let $\mathcal{A}=\Delta$ be the $d$-simplex of discrete probability distributions
$q=(q(x_{1}),\dots,q(x_{d}))$
on a finite space $\mathcal{X}$ with $d$ elements. The log loss defined on $\mathcal{X}\times\mathcal{A}$ is given by
$l(x,q)=-\log q(x).$
It is well-known that the log loss is a so-called proper loss, justifying its use in probability estimation, since on average $\log q(X)$ is minimal at the distribution $q$ over which the average is taken. In particular, for any two probability distributions $p,q\in\mathcal{A}$, we have
$\mathbb{E}_{X\sim p}[-\log q(X)]\geq\mathbb{E}_{X\sim p}[-\log p(X)].$
Moreover, there is no remaining noise term in $H(X)$ for the log loss, as $\operatorname*{\arg\!\min}_{Q\,\sigma(X)\text{-meas.}}\mathbb{E}[-\log Q(X)]=\delta_{X}$, where $\delta_{x}$ denotes the Dirac measure centered at $x\in\mathcal{X}$. Note that here and in the optimization problems over $\sigma(Y)$-measurable random variables $A$ appearing in the expressions for $H(X|Y)$ and $I(X;Y)$, the optimization is over random probability measures $Q$. In total, we obtain
$H(X)$ $=\mathbb{E}[-\log\mathbb{P}_{X}]=H_{S}(X),$
$H(X|Y)$ $=\mathbb{E}[-\log\mathbb{P}_{X|Y}]=H_{S}(X|Y),$
$I(X;Y)$ $=\mathbb{E}[\log\mathbb{P}_{X|Y}-\log\mathbb{P}_{X}]=I_{S}(X;Y),$
$I(X;Y|Z)$ $=\mathbb{E}[\log\mathbb{P}_{X|Y,Z}-\log\mathbb{P}_{X|Z}]=I_{S}(X;Y|Z)\,$,
where we suppress the arguments of the conditional distributions inside the expectations for brevity. This recovers non-conditional and conditional Shannon entropy and Shannon information.
For the continuous case, let $\lambda$ be a fixed $\sigma$-finite measure on a measurable space $\mathcal{X}$, e.g., the Lebesgue measure on the real line $\mathbb{R}$. If we let $\mathcal{A}$ represent all probability densities $f$ on $\mathcal{X}$ with respect to $\lambda$, then the log loss on $\mathcal{X}\times\mathcal{A}$ is defined by
$l(x,f)=-\log f(x)$
for $\lambda$-almost every $x\in\mathcal{X}$ and $f\in\mathcal{A}$. Note that each $f\in\mathcal{A}$ corresponds to a probability measure on $\mathcal{X}$ that is absolutely continuous with respect to $\lambda$ and uniquely determined by $f$ up to a set of measure-zero. As can be seen by Jensen’s inequality, the log loss is proper also in the continuous case [35], in particular, $\inf_{G\,\{\emptyset,\Omega\}\text{-meas.}}\mathbb{E}_{f}[-\log G(X)]=\mathbb{E}_{f}[-\log f(X)]$, where we write $\mathbb{E}_{f}$ for the expectation with respect to the measure $fd\lambda$. This expression is known as the continuous or differential entropy $h(X)$, which is not a direct generalization of Shannon entropy. In particular, when taking the limit of smaller and smaller discretizations of the real line, Shannon entropy does not converge to $h(X)$. It becomes infinitely large, corresponding to the fact that one would have to use infinitely long binary sequences in order to encode each X=xX\,{=}\,x precisely. The minimization over $\sigma(X)$-measurable $A$ in our definition of $H(X)$ leads to $H(X)=\infty$ simply because $\inf_{G\,\sigma(X)\text{-meas.}}\mathbb{E}_{f}[-\log G(X)]=$-$\infty$, as we can find a sequence $G_{n}$ of density-valued functions of $x$ that take larger and larger values at $x$ (so that $G_{n}\lambda$ are approximations to the Dirac measure on $\mathcal{X}$ centered at $x$). More precisely, we can choose $\sigma(X)$-measurable $G_{n}^{x}$ to be a Gaussian density with mean $x$ whose variance converges to zero as $n\to\infty$, for example $G_{n}^{x}(\xi)=\frac{n}{\sqrt{\pi}}e^{-n^{2}(\xi-x)^{2}}$. We have
$\inf_{A\,\sigma(X)\text{-meas.}}\mathbb{E}_{f}[l(X,A)]\leq\mathbb{E}_{f}[l(X,G_{n}^{X})]=-\log\frac{n}{\sqrt{\pi}}\,$,
in particular, $\inf_{A\,\sigma(X)\text{-meas.}}\mathbb{E}_{f}[l(X,A)]=$-$\infty$.
While entropy is infinite, for (unconditional and conditional) information in $Y$ about $X$, where uncertainty is not reduced fully but only to the knowledge level given by $Y$, we obtain results analogous to the discrete case. In summary, we have
$H(X)$ $=\infty,$
$H(X|Y)$ $=\infty,$
$I(X;Y)$ $=\mathbb{E}[D_{\mathrm{KL}}(f_{X|Y}\|f_{X})],$
$I(X;Y|Z)$ $=\mathbb{E}[D_{\mathrm{KL}}(f_{X|Y,Z}\|f_{X|Z})]\,$,
given that the corresponding (conditional) densities exist and are integrable almost everywhere. Reflecting the fact that we can store an arbitrary amount of information in a continuous random variable, $H(X)$ and $H(X|Y)$ are infinite. In particular, we did not introduce any constraint on the capability of observing the values of $X$, as one would do in practice, where one might, for example, only allow a maximum length of the floating point representations.
Example 3 (Hyvärinen loss).
Utilizing both the square error and the log loss, the Hyvärinen objective
$J(g)=\mathbb{E}_{f}\big[\|\nabla\log g(X)-\nabla\log f(X)\|^{2}\big]$ (29)
corresponds to matching the so-called scores $\nabla\log g$ and $\nabla\log f$ of densities $g$ and $f$ satisfying appropriate regularity conditions. Note that $J(g)$ is unrelated to Fisher information, which has a similar functional form, but takes gradients with respect to density parameters rather than the random variables over which the densities are defined. Introduced originally as a loss function for statistical inference [42], the Hyvärinen loss has influenced various approaches to score matching in the machine learning literature, in particular for training energy-based models (EBMs), for example for image denoising and super-resolution [46], and generative modeling in general [80, 81]. One of its appealing properties is the invariance of $\nabla\log f$ under multiplicative rescaling of $f$, enabling the use of unnormalized density approximations in the objective function $J$ without having to determine the normalization constants. Moreover, as can be shown by using integration by parts [42], the minimizer of (29) can be rewritten as $\operatorname*{\arg\!\min}_{g}\mathbb{E}_{f}[l(X,g)]$, with the Hyvärinen loss
$l(X,g)=\frac{1}{2}\|\nabla\log g\|^{2}+\Delta\log g\,.$ (30)
Similarly to the log loss, uncertainty reduction from no or some knowledge to full knowledge about $X$ is infinite, because $l(x,g)$ can be made arbitrarily large for each $x$ separately, by using test functions $g$ with increasing curvature. More precisely, we can shift $G_{n}^{x}$ defined in the previous example so that the Laplacian diverges. For $\tilde{G}_{n}^{x}(\xi)=G_{n}^{x}(\xi+\frac{1}{n^{2}})$, we have
$\nabla_{\xi}\log\tilde{G}_{n}^{x}(\xi)\Big|_{\xi=x}=-2n^{2}\Big(\xi-x+\frac{1}{n^{2}}\Big)\Big|_{\xi=x}=-2$
and
$\Delta_{\xi}\log\tilde{G}_{n}^{x}(\xi)\Big|_{\xi=x}\propto-n^{2},$
so that also $\inf_{A\,\sigma(X)\text{-meas.}}\mathbb{E}_{f}[l(X,A)]=$-$\infty$. In summary,
$H(X)$ $=\infty,$
$H(X|Y)$ $=\infty,$
$I(X;Y)$ $=\mathbb{E}[\|\nabla\log f_{X|Y}-\nabla\log f_{X}\|^{2}],$
$I(X;Y|Z)$ $=\mathbb{E}[\|\nabla\log f_{X|Y,Z}-\nabla\log f_{X|Z})\|^{2}]\,$,
if the corresponding densities have the required regularity and integrability properties. Hence, under the Hyvärinen loss, information $I(X;Y)$ equals the expected reduction in squared error when estimating the score function of $X$. Intuitively, it quantifies how much observing $Y$ improves our knowledge of the local geometry of the log-density of $X$, in the sense that it captures the change in direction of the steepest increase of $\log$-density of $X$ from no knowledge to the knowledge of $Y$.
As we have seen, theoretically one can store an arbitrary amount of information with respect to the log loss and the Hyvärinen loss in a continuous random variable. In both cases, the reason is that the loss function depends pointwise on the density or its derivatives and the action space consists of arbitrary densities, corresponding to perfectly fine-grained control. For example, if one were to restrict the action space to Gaussian distributions with an upper bound on the standard deviations, then we could not exploit this pointwise dependence to blow up the loss when $X$ is known perfectly. Despite the theoretical limits of $H(X)$ and $H(X|Y)$ being infinite, it is also clear from our definitions that $I(X;Y)$ and $I(X;Y|Z)$ can be finite, because they are not defined through entropy differences of the form H(X)H(X|Y)H(X)-H(X|Y) (where infinities would have to ’cancel out’) but through uncertainty reduction from no or partial information to partial information.
Example 4 (Bregman losses).
Let $\mathcal{X}$ be (a convex subset of) a Euclidean space. Bregman losses (or Bregman divergences) [14, 20] are a large class of loss functions of the form
$l(x,a)=d_{\phi}(x,a)\coloneqq\phi(x)-\phi(a)-\langle\nabla\phi(a),x-a\rangle\,$, (31)
where, $\phi:\mathcal{X}\to\mathbb{R}$ is a given convex and differentiable function on $\mathcal{X}$, and $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product on $\mathcal{X}$. Bregman losses can also be defined on function spaces where ordinary derivatives are not defined, for example by making use of Fréchet derivatives [26], or by using so-called subgradients $\phi^{\prime}$ of the convex functions $\phi$ [61], which are (generally non-unique) mappings $\phi^{\prime}$ from the underlying space to its algebraic dual such that $\phi(x)\geq\langle\phi^{\prime}(y),x-y\rangle+\phi(y)$, where $\langle\cdot,\cdot\rangle$ denotes the corresponding dual pairing. For example, if the underlying space consists of probability measures $p$, then one can use integrable functions as dual and the dual pairing is given by $\langle f,p\rangle=\int fdp$. The inner product in (31) is then replaced by the dual pairing and the gradient by a subgradient, see Equation (34) in Appendix -C.
Many loss functions that appear in the literature can be represented as Bregman divergences, see for example Table 1 in [6] or the examples in [61]. The basic example of a Bregman loss is the square error, which is generated by $\phi(x)=\|x\|^{2}$. Bregman divergences are tied to conditional expectations, as they are the only loss functions [5] that fulfill
$\operatorname*{\arg\!\min}_{A\,\mathcal{F}\text{-measurable}}\,\mathbb{E}[d_{\phi}(X,A)]=\mathbb{E}[X|\mathcal{F}]\,$, (32)
i.e., $\min_{A\,\mathcal{F}\text{-meas.}}\,\mathbb{E}[d_{\phi}(X,A)]=\mathbb{E}[d_{\phi}(X,\mathbb{E}[X|\mathcal{F}])]$ (see Appendix -C for a general discussion). Note that, by using the definition of $d_{\phi}$, this minimum can be further simplified to
$\mathbb{E}[d_{\phi}(X,\mathbb{E}[X|\mathcal{F}])]=\mathbb{E}[\phi(X)]-\mathbb{E}(\phi(\mathbb{E}[X|\mathcal{F}]))\,$,
in particular, the minimum is independent of the chosen subgradient in the functional case. We obtain
$H(X)$ $=\mathbb{E}[\phi(X)]-\phi(\mathbb{E}[X])\eqqcolon H_{\phi}(X),$
$H(X|Y)$ $=\mathbb{E}[\phi(X)-\phi(\mathbb{E}[X|Y])]=\mathbb{E}[H_{\phi}(X|Y)],$
$I(X;Y)$ $=\mathbb{E}[\phi(\mathbb{E}[X|Y])]-\phi(\mathbb{E}[X])=H_{\phi}(\mathbb{E}[X|Y]),$
$I(X;Y|Z)$ $=\mathbb{E}_{Z}[H_{\phi}(\mathbb{E}[X|Y,Z])],$
generalizing the corresponding expressions for the square error discussed in Example 1. In the expression for $I(X;Y|Z)$, the application of $H_{\phi}$ is only over the variable $Y$, so that the resulting expression still depends on $Z$. In the literature, $H_{\phi}$ is known as Bregman information [4] or Jensen gap, since it is the difference between the two sides of Jensen’s inequality. Bregman information has been shown recently to represent the variance term in a general version [34] of the well-known bias-variance trade-off [29]. In particular, it has been applied to quantify the prediction uncertainty of neural networks trained with the corresponding scoring rule [71].
Besides the square error, another well-known Bregman loss is the KL divergence, $l(p,q)=D_{\mathrm{KL}}(p\|q)$ defined on $\Delta^{2}$, where $\Delta$ is the $d$-simplex of probability distributions on a finite set with $d$ elements. In general, the random elements $X$ and $A$ in this case are random distributions. For example, if $X=\mathbb{P}_{Y|Z}$, which is a $\sigma(Z)$-measurable random distribution on the finite target space $\mathcal{Y}$, and $A=Q_{Y}$ is $\{\emptyset,\Omega\}$-measurable, i.e., it is a single distribution $q=(q(y_{1}),\dots,q(y_{d}))$, then by property (32),
$\operatorname*{\arg\!\min}_{Q_{Y}\,\{\emptyset,\Omega\}\text{-meas.}}\mathbb{E}[D_{\mathrm{KL}}(\mathbb{P}_{Y|Z}\|Q_{Y})]=\mathbb{E}_{Z}[\mathbb{P}_{Y|Z}(Y|Z)]=\mathbb{P}_{Y}\,$,
which plays an important role in many areas closely related to information theory [17], such as the rate-distortion literature [67], the information bottleneck method [76], and information-theoretic bounded rationality [30], to name a few. The minimum is the Shannon information, in particular, we have
$H_{\mathrm{KL}}(\mathbb{P}_{Y|Z})=\min_{Q_{Y}\,\{\emptyset,\Omega\}\text{-meas.}}\mathbb{E}[D_{\mathrm{KL}}(\mathbb{P}_{Y|Z}\|Q_{Y})]=I_{S}(Y;Z),$
which means that the total information in $\sigma(Z)$ about $\mathbb{P}_{Y|Z}$ in terms of the KL loss equals the information in $\sigma(Z)$ about $Y$ in terms of the log loss. The KL divergence thus can be seen as measuring information in the space of probability distributions in such a way that it matches the corresponding information with respect to the log loss.
As we discuss in the following, this relation between the log loss and KL divergence generalizes to loss functions for which the corresponding infima are attained.
Example 5 (Losses with Bayes acts).
As shown in [38], Shannon information $I_{S}(X;Y)$ quantifies the cost of transforming $\mathbb{P}_{X}$ into $\mathbb{P}_{X|Y}$ in terms of the average length of binary strings transmitted from a sender that knows $Y$ to a receiver that wants to generate samples from $\mathbb{P}_{X|Y}$. The interpretation of information as the cost of transforming distributions is not limited to Shannon information, as mentioned in Section III-B, where we have seen that the risks $R(Y)$ depend on the random variables $Y$ only through their conditional probability measures $\mathbb{P}_{X|Y}$. For loss functions for which so-called Bayes acts [25] exist, this view becomes explicit, because entropy and information can be directly expressed as divergences between distributions, as we discuss in the following.
A simple example of such a loss function in the discrete case is the zero-one loss [35], $l_{01}(x,a)=1-\delta_{x,a}$, where $\mathcal{A}=\mathcal{X}$ and $\delta_{i,j}$ denotes the Kronecker-Delta. Since for any $a\in\mathcal{A}$, we have $\mathbb{E}[l_{01}(X,a)]=1-\mathbb{P}_{X}(a)$, it follows for any random variable $Y$ that
$\inf_{A\ \sigma(Y)\text{-meas.}}\mathbb{E}[l_{01}(X,A)]=1-\mathbb{E}[\mathbb{P}_{X|Y}(a(Y))],$
where $a(y)\coloneqq\mathrm{argmax}_{a}\mathbb{P}_{X|Y=y}(a)$ denotes the Bayes act with respect to $\mathbb{P}_{X|Y}$.
In general, for a probability measure $q$ on $\mathcal{X}$, a Bayes act $a_{q}$ is an optimal action with respect to the expected value of $l$, i.e., $a_{q}\in\operatorname*{\arg\!\min}_{a\in\mathcal{A}}\mathbb{E}_{q}[l(X,a)]$. As we have seen in Examples 2 and 3, not all loss functions and sub-$\sigma$-algebras have corresponding Bayes acts (see, e.g., [15] for sufficient conditions). If there are multiple minima, following [35], we assume that one fixes a specific Bayes act $a_{q}$ for each $q$. If the corresponding Bayes acts exist, we find
$H(X)$ $=\mathbb{E}[D(\delta_{X}\|\mathbb{P}_{X})],$
$H(X|Y)$ $=\mathbb{E}[D(\delta_{X}\|\mathbb{P}_{X|Y})]\,$,
$I(X;Y)$ $=\mathbb{E}[D(\mathbb{P}_{X|Y}\|\mathbb{P}_{X})],$
$I(X;Y|Z)$ $=\mathbb{E}[D(\mathbb{P}_{X|Y,Z}\|\mathbb{P}_{X|Y})],$
where
$D(p\|q)\coloneqq\mathbb{E}_{p}[l(X,a_{q})-l(X,a_{p})]$ (33)
is the divergence [60, 21] between two measures $q$ and $p$ on $\mathcal{X}$ induced by the proper scoring rule $S(x,q)\coloneqq l(x,a_{q})$. Note that, if $l$ is itself a proper scoring rule, then $a_{q}=q$ and $D$ reduces to the divergence of $l$. For example, in the case of the log loss for discrete random variables, we have $S(x,q)=-\log q(X)$ and $D(p\|q)=D_{\mathrm{KL}}(p\|q)$. More examples of a proper scoring rules include the Tsallis score [78] and Brier score [16], for which $D$ is the Tsallis divergence and $L^{2}$-divergence of densities, respectively [21]. Applications of Tsallis and Brier scores include weather forecasting, statistical mechanics, and calibration of probabilistic classifiers.
The above results show that, for our knowledge-lattice-based definitions of entropy and information, we find the same relationships to $D$ as Shannon entropy and information have to $D_{\mathrm{KL}}$. Moreover, from the Bregman representation [56, 39, 31] of proper scoring rules, it follows that $D$ is also a Bregman divergence, just like KL divergence. In particular, we have
$\min_{Q_{Y}\,\{\emptyset,\Omega\}\text{-meas.}}\mathbb{E}[D(\mathbb{P}_{Y|Z}\|Q_{Y})]=\mathbb{E}[D(\mathbb{P}_{Y|Z}\|\mathbb{P}_{Y})],$
$\min_{Q_{Y}\,\sigma(Z)\text{-meas.}}\mathbb{E}[D(\mathbb{P}_{Y|Z}\|Q_{Y})]=\mathbb{E}[D(\mathbb{P}_{Y|Z}\|\mathbb{P}_{Y|Z})]=0\,$,
which means that fully reducing the uncertainty about $\mathbb{P}_{Y|Z}$ from $\{\emptyset,\Omega\}$ to $\sigma(Z)$ measured under $D$, i.e., the entropy $H(\mathbb{P}_{Y|Z})$ with respect to $D$, is the same as $I(Y;Z)$, the reduction in uncertainty about $Y$ from $\{\emptyset,\Omega\}$ to $\sigma(Z)$ measured with respect to $l$, generalizing the corresponding result for the KL divergence (cf. previous example).
For simplicity and in agreement with the literature on proper scoring rules, we have considered $D(p\|q)$ only for losses where both Bayes acts exist, $a_{q}$ and $a_{p}$. However, we could extend the above results to a more general definition of $D(p\|q)$ through $\sup_{a}\mathbb{E}_{p}[l(X,a_{q})-l(X,a)]$, only requiring $a_{q}$ to exist.
Figure 9: A possible sequence of decreasing risks that decompose $I(X;(X_{1},X_{2},X_{3}))$ into 32 non-negative terms. We write $\pi_{i}$ to denote $\pi_{X_{i}}=\pi(\mathbb{P}_{X|X_{i}})$ and $\pi_{i,j}$ for $\pi_{X_{i},X_{j}}=\pi(\mathbb{P}_{X|X_{i},X_{j}})$.
-C Background on Bregman divergences
Consider the general shape of a (functional or non-functional) Bregman divergence
$d_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\phi^{\prime}(y),x-y\rangle$ (34)
for a convex function $\phi$, a subgradient $\phi^{\prime}$ (gradient in the non-functional case), and the dual pairing $\langle\cdot,\cdot\rangle$ between the underlying space $\mathcal{X}$ and its algebraic dual $\mathcal{X}^{\prime}$ (see Example 4 in Section II). A simple calculation shows
$d_{\phi}(x,y)-d_{\phi}(x,z)=d_{\phi}(z,y)+\langle\phi^{\prime}(z){-}\phi^{\prime}(y),x{-}z\rangle$ (35)
for arbitrary $x,y,z\in\mathcal{X}$, and thus, for $\mathcal{F}$-measurable random variables $Y$, we find
$\mathbb{E}[d_{\phi}(X,Y){-}d_{\phi}(X,\mathbb{E}[X|\mathcal{F}])|\mathcal{F}]=d_{\phi}(\mathbb{E}[X|\mathcal{F}],Y),$ (36)
if the dual pairing satisfies $\mathbb{E}[\langle\alpha,X\rangle]=\langle\alpha,\mathbb{E}[X]\rangle$ for any constant $\alpha\in\mathcal{X}^{\prime}$, which holds under fairly general conditions, see e.g. [11]. More precisely, (36) follows from (35) if
$\mathbb{E}[\langle\phi^{\prime}(Z),X\rangle|\mathcal{F}]=\langle\phi^{\prime}(Z),\mathbb{E}[X|\mathcal{F}]\rangle,$ (37)
whenever $Z$ is $\mathcal{F}$-measurable. For example, for non-functional Bregman divergences, where the dual pairing is a Euclidean inner product, (37) follows from $\mathbb{E}[YX|\mathcal{F}]=Y\,\mathbb{E}[X|\mathcal{F}]$ whenever $Y$ is $\mathcal{F}$-measurable, which is used in [5] to derive (32) from (36). In [34], the same calculation is used to show (36) with $\sigma=\{\emptyset,\Omega\}$ for functional Bregman divergences on random measures.
Equation (36) shows that the Bregman divergences from $X$ to any $\mathcal{F}$-measurable $Y$ can be decomposed into two terms consisting of the divergences from $X$ to $\mathbb{E}[X|\mathcal{F}]$ and $\mathbb{E}[X|\mathcal{F}]$ to $Y$,
$\mathbb{E}[d_{\phi}(X,Y)]=\mathbb{E}[d_{\phi}(X,\mathbb{E}[X|\mathcal{F}])]+\mathbb{E}[d_{\phi}(\mathbb{E}[X|\mathcal{F}],Y)],$
a general version of the Pythagorean theorem [20]. In particular, Bregman divergences (uniquely) keep the property of the square error, by which the divergence from $X$ to $Y$ can be written as the divergence of $X$ to its orthogonal projection $\mathbb{E}[X|\mathcal{F}]$ plus the divergence of the projection to the $\mathcal{F}$-measurable $Y$.
-D Proof of Proposition 3
The decomposition is a telescope sum and thus follows from the chain rule, once we have seen that
$R(\pi_{Y,Z})\leq R(\pi_{Y}\vee\pi_{Z}),$
because the other risks are trivially non-increasing since the corresponding partitions are even an ordered chain with respect to refinement, i.e.,
$\bot\ \preceq\ \pi_{Y}\wedge\pi_{Z}\ \preceq\ \pi_{Y}\ \preceq\ \pi_{Y}\vee\pi_{Z}\,$,
and the same with $\pi_{Z}$ instead of $\pi_{Y}$. Hence, it suffices to show that synergistic information is non-negative. First of all, by Proposition 2, $R(\pi_{Y,Z})=R(Y,Z)$. It is therefore enough to prove that $R(Y,Z)\leq R(\pi_{Y}\vee\pi_{Z})$. For the latter, note that $\mathbb{P}_{X|Y}$ is $\sigma(Y)$-measurable and $\mathbb{P}_{X|Z}$ is $\sigma(Z)$ measurable. In particular, $\sigma(\mathbb{P}_{X|Y},\mathbb{P}_{X|Z})\subset\sigma(Y,Z)$, and so the inequality follows from $\sigma(\mathbb{P}_{X|Y},\mathbb{P}_{X|Z})=\sigma(\pi_{Y}\vee\pi_{Z})$.
-E Partial information decompositions for three sources
Figure 9 illustrates one possible sequence of partitions of $\Omega$ along a path from $\bot$ to $\pi_{1,2,3}$ with decreasing risks, inducing a decomposition of $I(X;(X_{1},X_{2},X_{3}))$ through the chain rule in Proposition 1 $(v)$. The shown risk sequence corresponds to a decomposition with 32 terms, which is more fine-grained than the typical assumption of 18 informational terms commonly reported in the PID literature [33]. In contrast to previous PID formulations, these terms correspond to conditional risk reductions associated with successive transitions between knowledge states.
Similar to the case of two sources, in our experiments, we were able to find examples that show that each of the shown transitions is generally non-zero, which we omit for the sake of brevity.